ROVIBRATIONAL SPECTROSCOPY OF CO_{2}-NO

\underline{C. DUTTON}, S. EVANS, R. A. BEAUDET , Department of Chemistry, University of Southern California, Los Angeles, CA 90089-0482.

The absorption spectrum of the weakly bound cluster CO_{2}-NO has been observed by probing the v_{3} asymmetric stretch of CO_{2}. Clusters were formed by supersonic expansion of a mixture of NO:CO_{2} in a 7:1 ratio using He as the carrier gas. Etalon, reference and signal spectra were recorded simultaneously using a transient digitizer and a 386 PC. The spectrum was recorded between 2344.6 -2349.6 cm^{−1}. The structure of the dimer will be discussed.